OpenStudy (anonymous):

I think im the only one in the whole world that is clueless about functions Let f (x) = 2x +1 and g(x) = x 3 . If x = 4, what is the value of g(f(x))

OpenStudy (akshay_budhkar):

i think its 9???

OpenStudy (anonymous):

what is g(x)? is it g(x) = x+3?

OpenStudy (anonymous):

or just g(x)=x?

OpenStudy (anonymous):

yeah g(x) = x

OpenStudy (anonymous):

you sure? what is that 3 doing there?

OpenStudy (akshay_budhkar):

yeah lol

OpenStudy (anonymous):

oh I didn't see that

OpenStudy (anonymous):

maybe it is waiting to become a 4

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

lol

OpenStudy (akshay_budhkar):

lol

OpenStudy (anonymous):

g(x)= x/3

OpenStudy (akshay_budhkar):

then ur answer is 3 lol

OpenStudy (anonymous):

ooooooooooooooooh

OpenStudy (anonymous):

\[f(4)=9\] \[g(9)=\frac{9}{3}=3\]